OpenStudy (anonymous):

please help me solve for x log2(7x + 1) = log2(2 - x)

OpenStudy (anonymous):

Assuming it is "base 2", you have 7x + 1 = 2 - x. Solve for x now.

OpenStudy (anonymous):

6x+1 ?

OpenStudy (anonymous):

8x +1

OpenStudy (anonymous):

7x + 1 = 2 - x 7x + x = 2 - 1 8x = 1 x = 1/8

OpenStudy (anonymous):

why a fraction? thanks

OpenStudy (anonymous):

If 8 times x is 1, x has to be a fraction and be equal to 1/8.

OpenStudy (anonymous):

thank you. This is still hard for me

OpenStudy (anonymous):

1/8 + 1/8 + 1/8 + 1/8 + 1/8 + 1/8 + 1/8 + 1/8 = 1